OpenStudy (anonymous):

Suppose that someone wants to lay a pipeline around the earth. Also suppose that the circumference of the earth is exactly 25,000 miles. The manufacturer made the pipeline exactly 20 feet too long. Still, it was proposed to put it in position, supported above the surface of the earth by posts of equal length. How high above the surface of the earth would the pipeline be? Remember C=2pir, You may want to use h for the height of the posts. Round answer to the nearest hundredth and label answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

height* of posts I mean, sorry.

OpenStudy (campbell_st):

this is an interesting question as it assumes the pipe has no height...that is diameter... If it has a diameter... the top if the pipe is higher than the bottom of the pipe. just a thought...

OpenStudy (radar):

Circumfurence of the pipe is 25,000 miles plus 20 ft. Need similar units here, so convert 20 ft to mile units =3.787878 10^-3 or .003787878 , C=25.000.003787878 C=2PiR R=C/2Pi, R=25000.003787878/2Pi =3978.874179 miles radius of earth is 25000=2Pir=3978.873577 Subtract getting 6.02 10^-4 milie or3.17856 feet for pole height. Rounding to nearest hundreth 3.18 ft poles.

OpenStudy (anonymous):

132,000,000 feet of pipe and they made it 20 feet too long? that's like what, a 1/10millionth of a percent error? Just wait for the sun to go down. The temperature will drop a few degrees and the metal will contract and you'll get a perfect fit. :-)

OpenStudy (radar):